OpenStudy (anonymous):

John has 15 more baseball cards than Michael. Together they have a total of 55 cards. How many cards does Michael have? Let M = the number of baseball cards Michael has.

OpenStudy (anonymous):

let x= the number of baseball cards michael has let 15+x= the number of baseball cards john has

OpenStudy (charlotte123):

welcome to openstudy :D M= J +15 M+J =55 Solve for J in the first equation m= j +15 m- 15 = j Substitute in second equation m + (m-15) = 55 m + m-15=55 2m - 15 = 55 2m = 70 m=35 :D In other words, Michael has 35 cards ^_^

OpenStudy (anonymous):

if doing it with one variable you have x+x+15=55 and solve for x

OpenStudy (anonymous):

but these are the answers listed look.. a ) M + (M + 15) = 55 b ) M + (M - 15) = 55 c )M - (M + 15) = 55 d )M - (M - 15) = 55

OpenStudy (charlotte123):

to give a medal for the best answer u see, just click on best response :D enjoy ur stay here, and if u need help on anything else, just PM me ^_^

OpenStudy (charlotte123):

It would be B :D Micheal has 35 cards, subsitute M for 35, and u will get 35 + (35 - 15) = 55, meaning 35 + (20) = 55, which is true :D

OpenStudy (anonymous):

which is the answer a,b,c,d ?

OpenStudy (charlotte123):

it is B @q123456789 ^_^

OpenStudy (anonymous):

x+x+15 = 55 2x+15=55 solve for x

OpenStudy (anonymous):

x+x+15 = 55 so where can you but brackets and leave this statement true?